OpenStudy (anonymous):

Please show all of your work Describe the graph of 3x - 4y = 12 (Include the following in your answer) a) Slope b) y-intercept c) x-intercept d) Direction of the line (up right, down right) Thanks

OpenStudy (anonymous):

a) rearrange ur given formula into the form y=mx+b y=(3x/4)-3 so the slope is m therefore (3/4)

OpenStudy (anonymous):

b)for the y intercept set x =0 3(0)-4y=12 y=-3

OpenStudy (anonymous):

c) x intercept set y=0 3x-4(0)=12 x=4

OpenStudy (anonymous):

the line is sloping up (increasing)

OpenStudy (anonymous):

start off with y = mx + b that's your equation with m being your slope with b being your y intercept so use algebra to rearrange the equation to that. once you have the y intercept set y to zero and solve for x if you need help let me know